OpenStudy (anonymous):

(sin(x+60)cos((x-60)/2))/(sin((x/2)+60)sinx)=0.5

OpenStudy (anonymous):

\[\frac{(\sin(x+60)\cos((x-60)/2))}{(\sin((x/2)+60)sinx)}=0.5\]

OpenStudy (anonymous):